Two parallel plates of an air dielectric capacitor are separated by 5 mm to obtain a capacitance of 1F. What should be the area of ​​each plate?
Answer #1

C = capacitance = 1F

separated by 5 mm, distance d = 5mm = 0.005 m

A = area of ​​each plate = ?

eo = 8.85 x 10 -12

By Capacitance formula-

C = eo * A /d

A = C * d / eo

= 1 * (0.005) / (8.85 x 10-12)

= 5.65 x 108 m2